graph the line y+3=1/3(x-5) in two points
Answer by ewatrrr(24785)

y+3=1/3(x-5)
y = (1/3)x - 5/3 - 9/3
y = (1/3)x - 14/3
P(0, -14/3)
y=0, (1/3)x = 14/3, x=14
P(14, 0)